OpenStudy (anonymous):

Find all solutions to the equation 4tan(theta) + sqrt3= tan(theta)

OpenStudy (zehanz):

You can subtract tan(theta) and sqrt(3) from both sides, then you will have a much simpler equation.

OpenStudy (anonymous):

Possible answers are: 30º + n180º 300 + n180º 60º + n180º 150º + n180º

OpenStudy (anonymous):

You mean so I have 4*tan(theta on one side then tan(theta) - sqrt(3) on the other side?

OpenStudy (zehanz):

No, on the left side you have 4tan(theta). Subtract tan(theta), then you have 3tan(theta) left. (It's gone on the right side now). THen subtact sqrt(3) from both sides. Now you have:\[3\tan (\theta)=-\sqrt{3}\]

OpenStudy (zehanz):

YOu could see it this way: the equation looks like: 4a+b=a Subtract a from both sides: 3a+b=0 Subtract b from both sides: 3a=-b

OpenStudy (anonymous):

what do i do from there?

OpenStudy (zehanz):

YOu need theta, so knowing what tan(theta) is would help. Therefore, divide by 3 on both sides.

OpenStudy (anonymous):

I wasnt given theta :/

OpenStudy (zehanz):

No, theta is to be determined, and that is what we are busy doing right now ;) \[\tan(\theta)=-\frac{ 1 }{ 3 }\sqrt{3}\]

OpenStudy (anonymous):

ohhh ! ok so should i plug that into the equatyion?

OpenStudy (zehanz):

No, you've simplified the equation, and now it has become:\[\tan(\theta)=-\frac{ 1 }{ 3 }\sqrt{3}\] This is a known value of tan, there is a nice angle that goes with it:\[\tan(30º)=\frac{ 1 }{ 3 }\sqrt{3}\]so\[\tan(-30º)=-\frac{ 1 }{ 3 }\sqrt{3}\] -30º is not in the list, but if you add 180º, you get 150º. and adding any number of times 180º is also right. So the answer is D (last one)